During the night of Oct 20/21, we for the first time ran top-up injections into both rings during the same shift. The 3 GeV storage ring was being topped up between 2.9 and 3.0 mA (green trace in the graph above, Amps on right axis), while the 1.5 GeV storage ring was topped up between
On September 30, we stored beam for the first time in the MAX IV 1.5 GeV storage ring. This was enabled by injecting into the storage ring (roughly 240 pC per shot net charge in TR1) while powering both 100 MHz RF cavities. Shortly after storing about 0.08 mA (as shown in the screen shot
Today beam was injected from the linac to the 1.5 GeV storage ring and for the first time several turns in the storage ring could be detected. The image shows the BPM sum signal from the third BPM in achromat 1. It clearly reveals three beam passages separated by 320 ns, the ring’s revolution period.
Summer shutdown 2016 has started. In the 3 GeV storage ring we will be installing three new IDs: the in-vacuum wiggler for BALDER, EPU53 for HIPPIE, and EPU48 for VERITAS. These are our first EPUs and they will be joining the two already operational in-vacuum undulators. In the 1.5 GeV storage ring we are installing
After being able to ramp up the beam current to around 130 mA, we can start to see the stabilizing effect of our three Landau cavities operating at the third harmonic (300 MHz). For currents above ~125 mA, the shunt impedance of these cavities becomes sufficient to approach flat-potential conditions when the cavities are properly tuned
Today we ran the 3 GeV storage ring in decaying beam mode (with slow orbit feedback running at 1 Hz) in order to deliver photons to the NanoMAX and BioMAX beamlines for beamline commissioning. NanoMAX was able to get photons all the way to their monochromator. While we have been using dipole radiation today, tomorrow
We have started to close the gap of our first in-vacuum undulator IVU18 from Hitachi at the BioMAX beamline. The minimum gap setting we achieved today was 7 mm. With the gap closed we still see reasonable lifetime and orbit. Work will now start on compensating for the effects of this ID in order to
We have made progress in both the linac and the 3 GeV storage ring. Today we had 80 mA stored in the ring. An integrated beam lifetime of about 0.8 Ah has been recorded. Meanwhile we can also switch back and forth between photo-electron delivery to the SPF and top-up injection shots (electrons from the
Startup after the Feb/Mar shutdown is progressing. The linac has come up well and is delivering bunches all the way through the transfer line. The ring RF has come up, five cavities online. Now we just need to get beam stored in the ring.